HYDROCARBON RADICALS ISOLATED IN HELIUM NANODROPLETS

\underline{ALEXANDER M. MORRISON}, STEVEN D. FLYNN, TAO LIANG and GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602.

We report high resolution infrared spectra in the C-H stretch region for small hydrocarbon radicals isolated in helium nanodroplets. Stark Spectroscopy of the symmetric CH_{2} stretch of the ethyl radical provides the first experimentally measured dipole moment for this species.